OpenStudy (anonymous):

Need help setting up the Work/Force Integral: Crane lifts a bucket of sand originally weighing 125 lb at a constant rate. Sand is lost from the bucket at a constant rate of 0.5lb/ft. How much work is done in lifting the sand 90 ft?

OpenStudy (anonymous):

. . . . .

OpenStudy (anonymous):

work is being done against gravity only. therefore -ve of work done by gravity = work done by crane. hence, work done by gravity = \[\int\limits_{0}^{h} mgdx\] At any instant, At a height x above the ground, mass = 125 - 0.5x therefore work done = \[\int\limits_{0}^{h} [125 - 0.5x]gdx\] taking g as constant try integrating it now. remember the units. i hope this is correct.

OpenStudy (anonymous):

hermeezy, that is not the expression for force. i think.

OpenStudy (anonymous):

Here's what i did without setting up the integral: losing 0.5ft/lb in 90 ft is total of 45 lb of sand lost. 125-(45/2) = 102.5 <----average amount of sand in bucket W = fd = 102.5(90) = 9225 lb/ft Since i don't know how to set up a proper integral, this is the best i can come up with.

OpenStudy (anonymous):

why is average amount of sand in the bucket equal to your force? force is mass times acceleration.

OpenStudy (anonymous):

haha. ;)

OpenStudy (anonymous):

@siddhantsharan what is g in your equation? gravity ? 9.8?

OpenStudy (anonymous):

yep. gravitational acceleration it is. and when you find the answer, remember to convert it into ft/s^2. 9.8 is in si units.

OpenStudy (anonymous):

Is it any difference based on your equation if the question says "neglect the weight of the bucket" ? sorry should have include that in there.

OpenStudy (anonymous):

i neglected it in my equation. thought that if it was there you wouldve given it. :)

OpenStudy (anonymous):

you understood the solution?

OpenStudy (anonymous):

Still working it out, did you end up with an answer to your equation by any chance?

OpenStudy (anonymous):

umm \[-296805.15\]\[ft ^{2} lb s ^{-2}\]

OpenStudy (anonymous):

that ll be a positive, sorry.

OpenStudy (anonymous):

whoa...okay , i'll work on it. thanks!

OpenStudy (anonymous):

cheers. :)